# Probabilistic Fusion and Calibration of Neural Speaker Diarization Models

### v1

@@ Abstract
End-to-End Neural Diarization (EEND) systems produce frame-level probabilistic speaker activity estimates, yet since evaluation focuses primarily on Diarization Error Rate (DER), the reliability and calibration of these confidence scores have been largely neglected. When fusing multiple diarization systems, DOVER-Lap remains the only established approach, operating at the segment level with hard decisions. We propose working with continuous probability outputs, which enables more sophisticated calibration and fusion techniques that can leverage model uncertainty and complementary strengths across different architectures. This paper presents the first comprehensive framework for calibrating and fusing EEND models at the probability level. We investigate two output formulations—multilabel and powerset representations—and their impact on calibration and fusion effectiveness. Through extensive experiments on the CallHome two-speaker benchmark, we demonstrate that proper calibration provides substantial improvements even for individual models (up to 19% relative DER reduction), in some cases mitigating the absence of domain adaptation. We reveal that joint calibration in powerset space consistently outperforms independent per-speaker calibration, and that the Fuse-then-Calibrate ordering generally outperforms calibrating individual models before fusion while requiring calibration of only a single combined model. Our best configuration outperforms DOVER-Lap in terms of DER while providing reliable confidence estimates essential for downstream applications. This work proposes best practices for probability-level fusion of EEND systems and demonstrates the advantages of leveraging soft outputs over hard decisions.
@@ I Introduction
Speaker diarization identifies and temporally localizes individual speakers in multi-speaker audio recordings. Traditional modular diarization systems rely on assigning speech segments to speakers through a pipeline of speaker embedding extraction and clustering algorithms [6135543]. While effective, these systems produce hard segment-level assignments where the distance to cluster centroids can provide some measure of uncertainty. To handle speaker overlap, external modules are often incorporated into the pipeline. The emergence of End-to-End Neural Diarization (EEND) [fujita2019end] fundamentally changed this paradigm by directly predicting frame-level speaker activity probabilities for each time frame in an audio stream, enabling integrated modeling of overlapping speech and providing explicit confidence estimates at the frame level.
@@
While recent EEND models achieve remarkably low error rates in speaker activity detection, both the reliability of their confidence estimates and the fusion of multiple diarization systems remain largely unexplored. Despite the critical role that confidence estimates play in downstream applications and model fusion, research on calibration of neural diarization models is extremely limited, with only one recent study addressing this topic [plaquet24_interspeech]. This gap becomes particularly critical when combining predictions from multiple models, as each model may exhibit distinct confidence patterns and biases that can significantly degrade fusion quality without proper calibration. The rapid evolution of EEND architectures has produced a diverse ecosystem of models with varying strengths: some excel at capturing long-range temporal dependencies [10848627, eend-vector-clustering], others at handling overlapping speech [9688044], and still others at speaker discrimination in noisy conditions [10022924].
@@
This diversity naturally raises the question of whether these complementary capabilities can be effectively combined. Existing fusion approaches like DOVER [9004031] and DOVER-Lap [DoverLap21] operate on hard decisions at the segment level, limiting their ability to leverage the probabilistic nature of EEND outputs. We show that naive probabilistic model averaging is not able to effectively combine EEND models, and more advanced fusion techniques are required. We argue that this incompatibility arises from model miscalibration—that is, their confidence predictions cannot be reliably interpreted as actual class probabilities [Dawid, Brocker09, degroot83]. Modern neural networks (NNs), the backbone of EEND models, have been shown to produce overconfident predictions [guo2017calibration, Ovadia2019]: for instance, when a network assigns 0.8 confidence to its predictions, the actual accuracy may be substantially lower than 80%. Given that EEND systems provide frame-level probabilistic outputs, they are uniquely suited for calibration and fusion techniques that require continuous probability estimates—an approach not applicable to methods like DOVER-Lap that operate on hard segment-level decisions.
@@
This paper addresses these fundamental limitations towards a framework for calibrating and fusing neural diarization models. While post-hoc calibration techniques have been successfully applied in speaker recognition tasks [vanleeuwen2007introduction, 4291590], their application to speaker diarization presents unique challenges. Our key insight is that the multilabel nature of speaker diarization, where multiple speakers can be simultaneously active, requires rethinking traditional calibration and fusion approaches designed for single-label classification problems. Moreover, well-calibrated probabilities enable optimal decision-making under varying costs and priors through decision theory—an important avenue for future work.
@@
We investigate two output formulations for this task: the standard multilabel formulation that treats speakers independently [fujita2019end], and the powerset formulation that explicitly models one class for each of all speaker combinations [du2022speaker, plaquet2023powerset]. Hereafter we will refer to these as Mult (Multilabel) and Power (Powerset) formulations. These output representations lead to different fusion behaviors and calibration requirements that have not been previously explored in neural diarization.
@@
Through extensive experiments on the CallHome two-speaker benchmark [martin2000nist], we make several key contributions that represent, to our knowledge, the first systematic study of calibration and score-level fusion for EEND systems. First, we demonstrate that proper calibration provides substantial improvements even for individual EEND models, in some cases mitigating the absence of domain adaptation. Unlike prior work that operates on hard decisions, we show that calibrating frame-level probabilistic outputs enables more effective model combination. Second, we explore both unsupervised fusion methods and supervised approaches for combining EEND outputs, analyzing how each interacts with calibration strategies and domain adaptation. Third, we reveal that the choice between multilabel and powerset representations fundamentally impacts both calibration quality and fusion effectiveness, with certain methods benefiting from explicit modeling of speaker combinations while others performing similarly across both formulations. Importantly, we find that improvements in calibration quality (measured by Binary Cross-Entropy [ferrer2024evaluating, e20030208]) generally translate to corresponding improvements in diarization performance (measured by DER), demonstrating that better-calibrated confidence estimates support more accurate speaker discrimination. Finally, we establish best practices for score-level fusion of EEND speaker diarization systems, demonstrating that fusing first and then calibrating (Fuse-then-Calibrate) generally outperforms calibrating individual models before fusion, and that joint calibration in powerset space consistently outperforms independent per-speaker calibration.
@@ II-A End-to-End Neural Diarization
In this work, we utilize pre-trained EEND with Encoder-Decoder-based Attractors (EEND-EDA) [horiguchi20_interspeech] models with different input feature representations as the foundation for our calibration and fusion framework. Following the approach in [app15094842], we leverage multiple EEND-EDA models trained with diverse feature sets to exploit their complementary speaker information. These models serve as the base components for our calibration and fusion methodology.
@@ II-B1 Multilabel Formulation
where $p_{s}^{\text{Mult}}$ represents the probability of speaker $s$ being active. These probabilities can be collected in a vector $\mathbf{p}^{\text{Mult}}=(p_{1}^{\text{Mult}},\ldots,p_{S}^{\text{Mult}})\in[0,1]^{S}$ for convenience.
@@ II-C Post-hoc Calibration of Probabilistic Classifiers
In this work, we employ Platt scaling [platt1999probabilistic] and its multi-class extension [niko_LRs], which essentially applies logistic regression to the classifier outputs. The specific implementation of this regression varies depending on the calibration strategy, with detailed methods presented in the next section.
The general approach minimizes the cross-entropy loss over a calibration set with respect to learnable parameters. Let $\mathbf{z}\in\mathbb{R}^{K}$ denote the pre-softmax output vector, or logit vector, of a classifier over $K$ classes. The method is implemented as:
@@ III-B2 Supervised Fusion Method
In contrast to the unsupervised approaches, we investigate a data-driven fusion strategy using a metalearner—a logistic regression model that learns to combine system outputs. The metalearner learns optimal combination weights from labeled data, taking the concatenated logits from all models as input:
@@ V-A Calibration of Base Models
The results show that individual base models benefit from calibration, especially those without fine-tuning. For example, MFB improves from 10.381% to 8.397% DER (19.1% relative reduction), while BCE improves from 0.419 to 0.271. After calibration, non-fine-tuned models achieve performance comparable to fine-tuned models: calibrated MFB (8.397% DER) approaches fine-tuned uncalibrated MFB (8.236% DER). This effect becomes more pronounced when fusing predictions: the fusion of calibrated models provides a 7.8% relative improvement in DER over raw fusion for non-fine-tuned models (8.425% to 7.764%), but only 0.5% for fine-tuned ones (7.067% to 7.031%). Nonetheless, calibration consistently improves performance across all model configurations, with BCE improvements ranging from 0.025 (FT GeMAPS) to 0.229 (No FT ECAPA-TDNN).
@@ VI Conclusions
Our key findings establish critical best practices for neural diarization: (1) Powerset representations with joint calibration consistently outperform independent per-speaker calibration (8.26% vs 9.18% DER), highlighting the importance of modeling speaker dependencies explicitly; (2) The Fuse-then-Calibrate strategy achieves superior performance (6.543% DER, 5.2% relative improvement over DOVER-Lap) while requiring calibration of only a single combined model; (3) Dynamic Logits fusion demonstrates the best performance across experimental conditions, effectively combining models at the logit level before softmax transformation.
@@
Future work should extend this framework to scenarios with more speakers, explore alternative proper scoring rules that may better correlate with DER optimization, and investigate applications in downstream tasks such as speaker-attributed ASR. By demonstrating that probability-level techniques outperform hard-decision approaches while providing reliable confidence estimates, this work establishes the foundation for more effective neural diarization systems that fully exploit their probabilistic outputs.

### v2

@@ Abstract
End-to-End Neural Diarization (EEND) systems produce frame-level probabilistic speaker activity estimates, yet since evaluation focuses primarily on Diarization Error Rate (DER), the reliability and calibration of these confidence scores have been largely neglected. When fusing multiple diarization systems, DOVER-Lap remains the only established approach, operating at the segment level with hard decisions. We propose working with continuous probability outputs, which enables more sophisticated fusion and calibration techniques that can leverage model uncertainty and complementary strengths across different architectures. This paper presents the first comprehensive framework for calibrating and fusing EEND models at the probability level. We investigate two output formulations—multilabel and powerset representations—and their impact on calibration and fusion effectiveness. Through extensive experiments on the CallHome two-speaker benchmark, we demonstrate that proper calibration provides substantial improvements even for individual models (up to 19% relative DER reduction), in some cases mitigating the absence of domain adaptation. We reveal that joint calibration in powerset space consistently outperforms independent per-speaker calibration, that fusion substantially improves over individual models, and that the Fuse-then-Calibrate ordering generally outperforms both calibrating before fusion and uncalibrated fusion while requiring calibration of only a single combined model. Our best configuration outperforms DOVER-Lap in terms of DER while providing reliable confidence estimates essential for downstream applications. This work proposes best practices for probability-level fusion of EEND systems and demonstrates the advantages of leveraging soft outputs over hard decisions.
@@ I Introduction
Speaker diarization identifies and temporally localizes individual speakers in multi-speaker audio recordings. Traditional modular diarization systems rely on assigning speech segments to speakers through a pipeline of speaker embedding extraction and clustering algorithms [6135543]. While effective, these systems produce hard segment-level assignments where the distance to cluster centroids could provide some measure of uncertainty although it is not usually considered. To handle speaker overlap, external modules are often incorporated into the pipeline. The emergence of End-to-End Neural Diarization (EEND) [fujita2019end] fundamentally changed this paradigm by directly predicting frame-level speaker activity probabilities for each time frame in an audio stream, enabling integrated modeling of overlapping speech and providing explicit confidence estimates at the frame level.
@@
While recent EEND models achieve remarkably low error rates in speaker activity detection, both the reliability of their confidence estimates and the fusion of multiple diarization systems remain largely unexplored. Despite the critical role that confidence estimates play in downstream applications and model fusion, research on calibration of neural diarization models is extremely limited. The only prior work [plaquet24_interspeech] assesses calibration to select poorly calibrated samples for retraining, but does not explore post-hoc calibration techniques or model fusion. This gap becomes particularly critical when combining predictions from multiple models, as each model may exhibit distinct confidence patterns and biases that can significantly degrade fusion quality without proper calibration. This gap becomes particularly critical when combining predictions from multiple models, as each model may exhibit distinct confidence patterns and biases that can significantly degrade fusion quality without proper calibration. The rapid evolution of EEND architectures has produced a diverse ecosystem of models with varying strengths: some excel at capturing long-range temporal dependencies [10848627, eend-vector-clustering], others at handling overlapping speech [9688044], and still others at speaker discrimination in noisy conditions [10022924].
@@
This diversity naturally raises the question of whether these complementary capabilities can be effectively combined. Existing fusion approaches like DOVER [9004031] and DOVER-Lap [DoverLap21] operate on hard decisions at the segment level, limiting their ability to leverage the probabilistic nature of EEND outputs. We show that naive probabilistic model averaging is not able to effectively combine EEND models, and more advanced fusion techniques are required. We argue that this incompatibility arises from model miscalibration—that is, their confidence predictions cannot be reliably interpreted as actual class probabilities [Dawid, Brocker09, degroot83]. Modern neural networks (NNs), the backbone of EEND models, have shown to produce overconfident predictions [guo2017calibration, Ovadia2019]: for instance, when a network assigns 0.8 confidence to its predictions, the actual accuracy may be substantially lower than 80%. Given that EEND systems provide frame-level probabilistic outputs, they are uniquely suited for calibration and fusion techniques that require continuous probability estimates—an approach not applicable to methods like DOVER-Lap that operate on hard segment-level decisions.
@@
This paper addresses these fundamental limitations towards a framework for calibrating and fusing neural diarization models. While post-hoc calibration techniques have been successfully applied in speaker recognition tasks [vanleeuwen2007introduction, 4291590], their application to speaker diarization presents unique challenges. Our key insight is that the multilabel nature of speaker diarization, where multiple speakers can be simultaneously active in a time-frame, requires rethinking traditional calibration and fusion approaches designed for single-label classification problems. Moreover, well-calibrated probabilities enable optimal decision-making under varying costs and priors through decision theory—an important avenue for future work.
@@
We investigate two output formulations for EEND speaker diarization: the standard multilabel formulation that treats speakers independently [fujita2019end], and the powerset formulation that explicitly models one class for each of all speaker combinations [du2022speaker, plaquet2023powerset]. Hereafter we will refer to these as Mult (Multilabel) and Power (Powerset) formulations. These output representations lead to different fusion behaviors and calibration requirements that have not been previously explored in neural diarization.
@@
Through extensive experiments on the CallHome two-speaker benchmark [martin2000nist], we make several key contributions that represent, to our knowledge, the first systematic study of calibration and score-level fusion for EEND systems. First, we demonstrate that proper calibration provides substantial improvements even for individual EEND models, in some cases mitigating the absence of domain adaptation. Unlike prior work that operates on hard decisions, we show that calibrating frame-level probabilistic outputs enables more effective model combination. Second, we explore both unsupervised fusion methods and supervised approaches for combining EEND outputs, analyzing how each interacts with calibration strategies and domain adaptation. Third, we reveal that the choice between multilabel and powerset representations fundamentally impacts both calibration quality and fusion effectiveness, with certain methods benefiting from explicit modeling of speaker combinations while others performing similarly across both formulations. Importantly, we find that improvements in calibration quality (measured by Binary Cross-Entropy [ferrer2024evaluating, e20030208]) generally translate to corresponding improvements in diarization performance (measured by Diarization Error Rate - DER), demonstrating that better-calibrated confidence estimates support more accurate multi-speaker segmentation. Finally, we propose best practices for score-level fusion of EEND speaker diarization systems, demonstrating that fusing first and then calibrating (Fuse-then-Calibrate) generally outperforms calibrating individual models before fusion, and that joint calibration in powerset space consistently outperforms independent per-speaker calibration.
@@ II-A End-to-End Neural Diarization
In this work, we utilize pre-trained EEND with Encoder-Decoder-based Attractors (EEND-EDA) [horiguchi20_interspeech] models with different input feature representations as the foundation for our calibration and fusion framework. Following our approach in [app15094842], we leverage multiple EEND-EDA models trained with diverse feature sets to exploit their complementary speaker information.
@@ II-B1 Multilabel Formulation
where $p_{s}^{\text{Mult}}$ represents the probability of speaker $s$ being active. Each $p_{s}^{\text{Mult}}$ is a probability from a one-vs-all binary classification problem, where $1-p_{s}^{\text{Mult}}$ represents the probability that speaker $s$ is not active. These probabilities can be collected in a vector $\mathbf{p}^{\text{Mult}}=(p_{1}^{\text{Mult}},\ldots,p_{S}^{\text{Mult}})\in[0,1]^{S}$ for convenience.
@@ II-C Post-hoc Calibration of Probabilistic Classifiers
In this work, we employ Platt scaling [platt1999probabilistic] and its multi-class extension [niko_LRs], which essentially applies flavours of logistic regression to the classifier outputs. The specific implementation of this regression varies depending on the calibration strategy, with detailed methods presented in the next section.
The general approach of Platt scaling, also known as matrix scaling, minimizes the cross-entropy loss over a calibration set with respect to learnable parameters. Let $\mathbf{z}\in\mathbb{R}^{K}$ denote the pre-softmax output vector, or logit vector, of a classifier over $K$ classes. The method is implemented as:
@@ III-B2 Supervised Fusion Method
In contrast to the unsupervised approaches, we investigate a data-driven fusion strategy using a so-called MetaLearner—a logistic regression model that learns to combine system outputs. The MetaLearner learns optimal combination weights from labeled data, taking the concatenated logits from all models as input:
@@ V-A Calibration of Base Models
The results show that individual base models benefit from calibration, especially those without fine-tuning. Since the calibration set coincides with the fine-tuning set, calibration provides not only better-calibrated probabilities but also a degree of domain adaptation. For example, MFB improves from 10.381% to 8.397% DER (19.1% relative reduction), while BCE improves from 0.419 to 0.271. After calibration, non-fine-tuned models achieve performance comparable to fine-tuned models: calibrated MFB (8.397% DER) approaches fine-tuned uncalibrated MFB (8.236% DER). This effect becomes more pronounced when fusing predictions: the fusion of calibrated models provides a 7.8% relative improvement in DER over raw fusion for non-fine-tuned models (8.425% to 7.764%), but only 0.5% for fine-tuned ones (7.067% to 7.031%). Nonetheless, calibration consistently improves performance across all model configurations, with BCE improvements ranging from 0.025 (FT GeMAPS) to 0.229 (No FT ECAPA-TDNN).
@@ VI Conclusions
Our key findings proposes critical best practices for neural diarization: (1) Powerset representations with joint calibration consistently outperform independent per-speaker calibration (8.26% vs 9.18% DER), highlighting the importance of modeling speaker dependencies explicitly; (2) The Fuse-then-Calibrate strategy achieves superior performance (6.543% DER, 5.2% relative improvement over DOVER-Lap) while requiring calibration of only a single combined model; (3) Dynamic Logits fusion demonstrates the best performance across experimental conditions, effectively combining models at the logit level before softmax transformation.
@@
Future work should extend this framework to scenarios with more speakers, explore alternative proper scoring rules that may better correlate with DER optimization, and investigate applications in downstream tasks such as speaker-attributed automatic speech recognition. By demonstrating that probability-level techniques outperform hard-decision approaches while providing reliable confidence estimates, this work establishes the foundation for more effective neural diarization systems that fully exploit their probabilistic outputs.